OpenStudy (anonymous):

There are (53)2 ⋅ 50 hens in a bird enclosure. What is the total number of hens in the enclosure? 0 55 56 530

OpenStudy (anonymous):

@jprose

OpenStudy (anonymous):

53 * 2 * 50 = 106 * 50 = 5300 is the problem written correctly ??

OpenStudy (anonymous):

so D